OpenStudy (anonymous):

An oxygen gas tank has a volume of 8 L and a pressure of 1500 N/m2. If the temperature decreases from 2500 K to 2000 K and its volume increases to 24 L, what is its final pressure? --------------------- 400 N/m2 625 N/m2 3600 N/m2 5625 N/m2 @Austin_Cochran

OpenStudy (anonymous):

im not so sure on this one. im sorry ill get someone.

OpenStudy (anonymous):

@McSpazzy

OpenStudy (anonymous):

or not.... umm..

OpenStudy (anonymous):

Im guessing C im pretty sure it is:)

OpenStudy (anonymous):

@Kryten will know

OpenStudy (anonymous):

ok first start with pv=nrt n=n n=pV/RT p1V1/RT1=p2V2/RT2 p1V1/T1=p2V2/T2 p1V1T2=p2V2T1 p2= p1V1T2/V2T1

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

i got 400 N/m2